DOPPLER-FREE STUDY OF THE Li_{2} MOLECULE: THE B^{1}Π_{u} STATE

N. BOULOUFA, \underline{P. CACCIANI}, and R. VETTER, Laboratoire Aimé Cotton (CNRS UPR3321), Campus d’Orsay, Bâtiment 505, 91405 Orsay, France.

We have studied the potential barrier in the B^{1}Π_{u} state of lithium dimers by use of a Doppler-free spectroscopy experiment in which a cw tunable laser beam crosses an effusive beam of lithium. Under these conditions, hundreds of wavenumbers have been measured to an accuracy of 10^{−3} cm^{−1}. From broadened profiles corresponding to P, Q and R transitions for v′=13, 14, 15 and 16, it is possible to describe the potential barrier of the B^{1}Π_{u} state, between 50 cm^{−1} and 150 cm^{−1} from its top. These results complement the ones obtained by use of Fourier transform spectroscopy with Doppler-broadened light sources^{a}. Theoretical interpretation is on the tracks.

aI. Russier, F. Martin, C. Linton, P. Crozet, A. J. Ross, R.Bacis and S. Churassy, J. Mol. Spectroscopy, 168, 39 (1994)